# Implementing feature binding through dendritic networks of a single neuron

### v1

@@ III-D Dendritic sublinearity implements the feature binding
We then consider a more difficult computing task of feature binding problem (FBP) [43]. For the input, we considered that the four input variables $\rm X_{i},i=1,2,3,4$ corresponded to four afferent neural ensembles of 500 neurons each in each branch, and that an ensemble was active, signaling $\rm X_{i}=1$. For the output, we considered a single somatic spike as the response to an input vector, such that if the neuron fires a spike then $\rm Y=1$, and no spike as $\rm Y=0$. Thus, to implement the FBP successfully, the cell should only fire spike in response to the input vectors $\rm X=[1,1,0,0],[0,0,1,1]$, and not to spike for the remaining four input vectors (Figure 8A). The dual feature binding problem (dFBP) is the opposite. The neuron should not fire spikes in response to the input vectors $\rm X=[1,1,0,0],[0,0,1,1]$, and only spike for the remaining four input vectors. Indeed, we found the neuron can implement FBP, for example, with a combination of branches 1, 3, 6, and 10 (Figure 8B, left), and dFBP with a combination of branches 1, 4, 11, and 18 (Figure 8B, right). Next, we want over 24 branches with a group of four branches for all input vectors and recorded the number of groups of branches for both FBP and dFBP tasks (Figure 8C) with asynchronous inputs. By varying the number of synaptic inputs, we found that implementing dFBP requires more inputs while FBP requires an appropriate number of synapses. As the number of synapses increases, the number of branches involved in the implementation of FBP increases. Compared to FBP, there are significantly fewer groups that can implement dFBP. Particularly, nearly all branches can be involved in FBP, while only half of the branches for dFBP over all 24 branches (Figure 8D). We analyzed how the branches involved in FBP and dFBP are distributed over the four input vectors, and found that $\rm X_{1}$ and $\rm X_{2}$ tend to sit in the one trunk and $\rm X_{3}$ and $\rm X_{4}$ on the other trunk (Figure 8E), due to that the two branches produce stronger sublinearity when they are on different main dendrites. For instance, the observed EPSP $\rm O(EPSP_{br1+br2})$ will be smaller than the expected $\rm E(EPSP_{br1+br2})$ when two branches on different main dendrites. Generally speaking, there is further sublinearity between different branch combinations. Such observations highly depend on the temporal input protocol. When synchronous input was used, dFBP was more easily implemented with fewer numbers of synaptic inputs (Supplemental Figure S5). In addition, the sublinearity index of each branch is calculated by the average of the sublinearity index of all dendrites within the branch (Supplemental Figure S6). From Figure 9A, it was observed that when inputs $\rm x_{1}$ and $\rm x_{2}$ are both distributed in branches with weak sublinearity, the neuron struggles to perform AND operations. However, the likelihood of achieving AND operations increases when at least one of the inputs is distributed in a branch with moderate sublinearity. Additionally, achieving effective feature binding problem (FBP) and double feature binding problem (dFBP) requires a balance of sublinearity within the four branches, emphasizing the need for a moderate balance of sublinearity to support these complex computations Figure 9B.
@@ IV Discussion
In the realm of single-neuron computation, the rational underpinning lies in the intricacies of nonlinear dendritic behavior. Here utilizing a Purkinje cell model with the detailed dendritic morphology, we revealed a distinctive feature of sublinear synaptic integration at different parts of dendrites. It is this inherent sublinearity within the dendritic architecture that empowers cerebellar PCs to implement binding of different features, potentially sending from parallel and climbing fibers. Our study sheds light on the computational prowess of PCs, emphasizing the pivotal role played by dendritic sublinearity in facilitating the realization of intricate neural computations.
@@ IV-A Sublinear dendritic integration in Purkinje cells
Neurons within the cerebellum manifest intricate nonlinear properties, a phenomenon deemed critical in orchestrating network functionality [44]. The dendrites of cerebellar interneurons intricately filter synaptic response times and amplitudes, engendering sublinear input-output relationships. Elucidating the transformation of temporally and spatially distributed inputs from granular cells constitutes a crucial pursuit in understanding their operational dynamics [21, 22]. Notably, the morphological attributes of neurons introduce a consequential dimension, wherein the amplitude and configuration of local EPSPs are subject to the nuanced intricacies of neuronal form [45].
In stark contrast, PCs present a more intricate dendritic morphology, necessitating the integration of a greater multitude of synaptic inputs. Our investigation reveals a distinctive property: PC dendrites manifest sublinear synaptic integration, with the degree of sublinearity accentuated in thin dendrites as opposed to their thicker counterparts. Remarkably, this position-dependent nonlinearity in synaptic integration is reminiscent of observations in CA1 pyramidal neurons [11]. A plausible rationale lies in the elevated input resistance of thin dendrites, mitigating the driving forces for synaptic current flow concerning their thicker counterparts. This underlines the PC’s ability to differentiate similar inputs based on the synaptic distribution position. Moreover, the integration of inputs from granule cells, strategically distributed along the spiny dendrites, undergoes effective filtering to preclude saturation.
@@ IV-B Dendrites sublinearity affects neuronal input-output dynamics
The intricate interplay between theory and experimentation has consistently stressed the critical role played by dendrites in shaping the input-output neuronal dynamics, conferring formidable computing capabilities upon individual neurons [55, 14, 15, 4]. The influence of nonlinear dendritic integration in actively modulating neuronal responses has been experimentally demonstrated across diverse cortical areas [56, 57, 58]. Theoretical insights suggest that dendritic superlinearity renders neurons sensitive to clustering, while sublinearity engenders a sensitivity to scattering [31, 43].
Our findings unveil a distinctive characteristic of PC dendrites, wherein pronounced sublinearity translates into global scatter sensitivity, particularly in response to asynchronous inputs. This outcome suggests a resilience of PCs against firing induced by clustered granule cell inputs, attributed to the abundance and sparsity of granule cells, thus avoiding the triggering of PC firing by a limited subset of granule cells. Remarkably, in the somatosensory cortex, features are localized within single dendrites [59], while synapses in the visual and auditory cortex’s layer 2/3 pyramidal neurons exhibit maximal responsiveness to inputs carrying distinct sensory features [60, 61]. Consequently, this global scatter strategy aligns with the neuron’s responsiveness to multi-feature discrete distributions. Furthermore, our investigation illuminates that weak sublinearity in dendrites favors the generation of bursts. Notably, the principal dendrite establishing a synaptic connection between climbing fibers and PC emerges as a region of weak sublinearity. This finding posits that climbing fiber inputs can more readily induce PCs to generate intricate complex spikes.
@@ IV-C Dendritic nonlinearity implements feature binding
In our study, we unveil the capability of PCs with sublinear dendritic characteristics, demonstrating their capacity to perform Boolean operations, primarily preferred AND operations. As the synaptic inputs increase, the execution of OR operations becomes more pronounced. Remarkably, the facilitation of AND operations is prominent when two inputs are strategically distributed on one of the two main trunk dendrites, however, the XOR operation is unachievable here, unlike human pyramidal cells [64]. Furthermore, the sublinear attributes of PC dendrites contribute to implementing feature binding problems, where both both asynchronous and synchronous input can realize dual feature binding problems. Intriguingly, during the implementation of the feature binding problem, inputs appear equally distributed among the main primary dendrites. This observation aligns with the notion that more primary dendrites are requisite for complex functions. Notably, human PCs [65, 66], characterized by more intricate dendritic structures and enhanced computing capabilities compared to rodents, typically exhibit 2-3 main trunks, in contrast to most rodents with only a single main trunk. The presence of multiple dendritic trunks is correlated with climbing fiber connections and long-term synaptic plasticity [67] and this innervation pattern can generate independent computational compartments within a single Purkinje cell [66]. Our findings further propose that branches between main dendrites boost sublinear synaptic integration, thereby augmenting the computational capability of a single neuron. Together with other recent studies, neurons endowed with nonlinear dendritic characteristics surpass their linear counterparts in computing more complex tasks [27].
IV-D Limitation
Ion channel currents play a crucial role in shaping the integrated properties of neurons, exemplified by the currents through $Na$ and $Ca^{2}$ channels. These currents serve to recompense synaptic currents lost through membrane resistance, thereby influencing the overall functionality of neurons [45]. Notably, the presence of dendritic spine $CaV_{2.3}$ voltage-sensitive calcium channels introduces a nonlinear regulatory mechanism for synaptic input within neurons [68]. PC dendrites exhibit an intricate distribution of ion channels, adding a layer of complexity to the nonlinear function [9, 69]. Moreover, Purkinje cells also receive inhibitory input from interneurons and significantly impact PC dynamics [70], which is essential for the precise timing and coordination of muscular activities controlled by the cerebellum [71, 72]. However, the effect of inhibitory input on the nonlinear synaptic integration ability of Purkinje cells has not been studied in depth here. However it is obvious that inhibition will further regulate the nonlinear synaptic integration ability of Purkinje cells, in particular, inhibitory strength depends on input location. Future research should investigate the impact of voltage-dependent ion channels and inhibitory input on dendritic nonlinearity to fully understand the complex mechanisms underpinning Purkinje cell function.
@@ IV-D Limitation
Figure S3: Related to Figure4. Left column: Diagrams of three synaptic input distribution patterns: Cluster Input: Synapses concentrated in a single dendrite. Local Scattered: Synapses distributed across a branch area. Global Scattered: Synapses spread across the entire dendritic tree. Middle column: Output spike frequency (Hz) as a function of input spike frequency (Hz) with different numbers of synapses (50: red, 100: green, 150: blue). Right column: Output spike frequency (Hz) as a function of input spike frequency (Hz) with different input rate (50 Hz: red, 100 Hz: green, 150 Hz: blue). Here is synchronous input.

### v2

@@ III-D Dendritic sublinearity implements the feature binding
We then consider a more difficult computing task of feature binding problem (FBP) [53]. For the input, we considered that the four input variables $X_{i},i=1,2,3,4$ corresponded to four afferent neural ensembles of 500 neurons each in each branch, and that an ensemble was active, signaling $X_{i}=1$. For the output, we considered a single somatic spike as the response to an input vector, such that if the neuron fires a spike then $Y=1$, and no spike as $Y=0$. Thus, to implement the FBP successfully, the cell should only fire spike in response to the input vectors $X=[1,1,0,0]$ and $[0,0,1,1]$, and not to spike for the remaining four input vectors (Figure 8A). The dual feature binding problem (dFBP) is the opposite. The neuron should not fire spikes in response to the input vectors $X=[1,1,0,0]$ and $[0,0,1,1]$, and only spike for the remaining four input vectors. Indeed, we found the neuron can implement FBP, for example, with a combination of branches 1, 3, 6, and 10 (Figure 8B, left), and dFBP with a combination of branches 1, 4, 11, and 18 (Figure 8B, right). Next, we want over 24 branches with a group of four branches for all input vectors and recorded the number of groups of branches for both FBP and dFBP tasks (Figure 8C) with asynchronous inputs. By varying the number of synaptic inputs, we found that implementing dFBP requires more inputs while FBP requires an appropriate number of synapses. As the number of synapses increases, the number of branches involved in the implementation of FBP increases. Compared to FBP, there are significantly fewer groups that can implement dFBP. Particularly, nearly all branches can be involved in FBP, while only half of the branches for dFBP over all 24 branches (Figure 8D). We analyzed how the branches involved in FBP and dFBP are distributed over the four input vectors, and found that $X_{1}$ and $X_{2}$ tend to sit in the one trunk and $X_{3}$ and $X_{4}$ on the other trunk (Figure 8E), due to that the two branches produce stronger sublinearity when they are on different main dendrites. For instance, the observed EPSP $O(EPSP_{br1+br2})$ will be smaller than the expected $E(EPSP_{br1+br2})$ when two branches on different main dendrites. Generally speaking, there is further sublinearity between different branch combinations. Such observations highly depend on the temporal input protocol. When synchronous input was used, dFBP was more easily implemented with fewer numbers of synaptic inputs (Figure S9). In addition, the sublinearity index of each branch is calculated by the average of the sublinearity index of all dendrites within the branch (Figure S10). It was observed that when inputs $X_{1}$ and $X_{2}$ are both distributed in branches with weak sublinearity, the neuron struggles to perform AND operations (Figure S11A). However, the likelihood of achieving AND operations increases when at least one of the inputs is distributed in a branch with moderate sublinearity. Additionally, achieving effective feature binding problem (FBP) and double feature binding problem (dFBP) requires a balance of sublinearity within the four branches, emphasizing the need for a moderate balance of sublinearity to support these complex computations (Figure S11B).
Moreover, nonlinear dendrites can achieve feature binding with time-dependent signals (Figure 9). The feature binding process is shown across three time steps ($t_{1},t_{2},t_{3}$), with four features distributed across dendritic branches (8, 7, 19, and 23). The membrane potential plots illustrate that spikes occur only when the third ($X_{3}$) and fourth ($X_{4}$) features are activated during the second ($t_{2}$) and third ($t_{3}$) time steps (Figure 9A). Importantly, changing the activation sequence fails to produce spikes, highlighting the order-dependent nature of this feature binding process. In this particular example, there are five True cases where FBP is successful, while only one Miss case shows the failure of FBP without spikes. This observation is seen from spike count across all 216 possible combinations of feature activations (Figure 9B). These results confirm that the dendritic nonlinearity allows precise temporal integration, enabling the neuron to selectively bind spatial and temporal features.
@@ IV Discussion
In the realm of single-neuron computation, the rational underpinning lies in the intricacies of nonlinear dendritic behavior. Here utilizing a Purkinje cell model with the detailed dendritic morphology, we revealed a distinctive feature of sublinear synaptic integration at different parts of dendrites. It is this inherent sublinearity within the dendritic architecture that empowers cerebellar PCs to implement binding of different features, potentially sending from parallel and climbing fibers. Our study sheds light on the computational prowess of PCs, emphasizing the pivotal role played by dendritic sublinearity in facilitating the realization of sophisticated neural computations.
@@ IV-A Sublinear dendritic integration in Purkinje cells
Neurons within the cerebellum manifest complex nonlinear properties, a phenomenon deemed critical in orchestrating network functionality [54]. The dendrites of cerebellar interneurons precisely filter synaptic response times and amplitudes, engendering sublinear input-output relationships. Elucidating the transformation of temporally and spatially distributed inputs from granule cells constitutes a crucial pursuit in understanding their operational dynamics [21, 22]. Notably, the morphological attributes of neurons introduce a consequential dimension, wherein the amplitude and configuration of local EPSPs are subject to the nuanced intricacies of neuronal form [55].
In contrast, PCs present a more complex dendritic morphology, necessitating the integration of a greater multitude of synaptic inputs. Our investigation reveals a distinctive property: PC dendrites manifest sublinear synaptic integration, with the degree of sublinearity accentuated in thin dendrites as opposed to their thicker counterparts. Remarkably, this position-dependent nonlinearity in synaptic integration is reminiscent of observations in CA1 pyramidal neurons [11]. A plausible rationale lies in the elevated input resistance of thin dendrites, mitigating the driving forces for synaptic current flow concerning their thicker counterparts. This underlines the PC’s ability to differentiate similar inputs based on the synaptic distribution position. Moreover, the integration of inputs from granule cells, strategically distributed along the spiny dendrites, undergoes effective filtering to preclude saturation.
@@ IV-B Dendrites sublinearity affects neuronal input-output dynamics
The intricate interplay between theory and experimentation has consistently stressed the critical role played by dendrites in shaping the input-output neuronal dynamics, conferring formidable computing capabilities upon individual neurons [65, 14, 15, 4]. The influence of nonlinear dendritic integration in actively modulating neuronal responses has been experimentally demonstrated across diverse cortical areas [66, 67, 68]. Theoretical insights suggest that dendritic superlinearity makes neurons sensitive to clustering inputs, while sublinearity enables sensitivity to scattering inputs [31, 53].
Remarkably, in the somatosensory cortex, features are localized within single dendrites [69], while synapses in the visual and auditory cortex’s layer 2/3 pyramidal neurons exhibit maximal responsiveness to inputs carrying distinct sensory features [70, 71]. Consequently, this global scatter strategy aligns with the neuron’s responsiveness to multi-feature discrete distributions. Furthermore, our investigation illuminates that weak sublinearity in dendrites favors the generation of bursts. Notably, the principal dendrite establishing a synaptic connection between climbing fibers and PC emerges as a region of weak sublinearity. This finding posits that climbing fiber inputs can more readily induce PCs to generate complex spikes (CSs). Simple spike (SS) activity and the wave-like pattern of CS activity have been shown to correlate with Zebrin expression [72]. Previous studies have shown that in Zebrin positive (Z+) PCs, CSs exhibit a stronger modulatory effect on SSs compared to Z- PCs [73]. Our findings reveal that high-frequency background stimulation suppresses CS generation, whereas low-frequency stimulation allows CS to occur more prominently. This is in line with the observation of the stronger CS modulation of SS in Z+ PCs shown in experiments [73], where the lower background activity favors CS occurrence under low-frequency conditions. We demonstrated that strong nonlinear dendritic responses consistently suppress CS generation, regardless of background frequency, offering fresh insights into the distinct computational roles of Z+ and Z- PCs in cerebellar processing.
@@ IV-C Dendritic nonlinearity implements feature binding
In our study, we unveil the capability of PCs with sublinear dendritic characteristics, demonstrating their capacity to perform Boolean operations, primarily preferred AND operations. As the synaptic inputs increase, the execution of OR operations becomes more pronounced. Remarkably, the facilitation of AND operations is prominent when two inputs are strategically distributed on one of the two main trunk dendrites, however, the XOR operation is unachievable here, unlike human pyramidal cells [76]. Furthermore, the sublinear attributes of PC dendrites contribute to implementing feature binding problems, where both asynchronous and synchronous input can realize dual feature binding problems. Intriguingly, during the implementation of the feature binding problem, inputs appear equally distributed among the main primary dendrites. This observation aligns with the notion that more primary dendrites are requisite for complex functions. Notably, human PCs [77, 78], characterized by more complex dendritic structures and enhanced computing capabilities compared to rodents, typically exhibit 2-3 main trunks, in contrast to most rodents with only a single main trunk. The presence of multiple dendritic trunks is correlated with climbing fiber connections and long-term synaptic plasticity [79] and this innervation pattern can generate independent computational compartments within a single Purkinje cell [78]. Our findings further propose that branches between main dendrites boost sublinear synaptic integration, thereby augmenting the computational capability of a single neuron. Together with other recent studies, neurons endowed with nonlinear dendritic characteristics surpass their linear counterparts in computing more complex tasks [27].
IV-D General functional implication of PC dendritic computation
The computational capabilities of PCs are fundamentally shaped by their complex dendritic properties, which underlie their various functional functions in the cerebellum. Studies have shown that dendritic morphology and synaptic plasticity enable PCs to perform pattern recognition and temporal sequence encoding [80, 81, 82]. The distribution of ion channels along dendrites facilitates multiplexed coding, allowing PCs to process multiple input streams simultaneously [50]. Interestingly, a similar demand for efficient multi-input integration appears in artificial systems. In computer vision, the ability to accurately process multiple input streams under affine transformations is essential for robust pose estimation and motion analysis [83, 84], reflecting a shared computational principle between biological and artificial domains. Nonlinear integration of dendritics further improves temporal discrimination, allowing advanced computations [85, 23]. Furthermore, increased dendritic complexity, as observed in human PCs, is associated with superior computational performance [86]. Dendritic nonlinearities enhance the computational capacity of neurons by facilitating multi-layered processing of information within individual cells.
Here we demonstrated a spatial gradient in sublinearity, with stronger suppression in thinner dendrites, which can improve PC computational capabilities by facilitating selective input filtering, localized dendritic processing, and efficient spatiotemporal integration across dendritic regions, as demonstrated in cerebellar interneurons [21]. Theoretical work by [53] proposed that sublinear dendritic integration enables scatter sensitivity, favoring distributed inputs. Interestingly, similar behavior has also been observed experimentally in cerebellar stellate cells, as reported in [87]. Our study confirms this in Purkinje cells, showing that dendritic sublinearity enforces a global scatter strategy, where only globally distributed inputs induce spiking. This outcome suggests a resilience of PCs against firing induced by clustered granule cell inputs, attributed to the abundance and sparsity of granule cells, thus avoiding the triggering of PC firing by a limited subset of granule cells. This selective filtering mechanism shares conceptual similarity with strategies developed in robust estimation problems in computer vision, where global consistency is prioritized over local agreement to resist noise and outliers [88].
Moreover, the FBP is particularly intriguing because it allows neurons to integrate multiple feature combinations into a cohesive representation [89, 90]. Analogously, in computer vision, binding multiple local features under geometric constraints is critical for achieving reliable correspondence estimation [91], a problem closely related to integrating distributed and heterogeneous inputs, as observed in biological neurons. In our study, by integrating temporal information into the feature binding problem, we demonstrated that PCs can selectively bind spatial and temporal features. This highlights the role of PCs in addressing complex problems that involve distributed and temporally ordered inputs [92, 93]. As described in [86], the increased dendritic complexity and computational capacity of human Purkinje cells support our findings, suggesting that the dendritic nonlinearity and spatiotemporal integration observed in rodents PCs could function even more efficiently in the more complex dendritic architecture of human PCs.
PCs also receive inhibitory input from interneurons that can significantly impact PC dynamics [94], which is essential for the precise timing and coordination of muscular activities controlled by the cerebellum [95, 96]. Inhibition could regulate the nonlinear synaptic integration ability of PCs. Such regulation may also depend on the input location of specific dendrites. Future research would investigate the impact of inhibitory input on dendritic nonlinearity to fully understand the complex mechanisms underpinning the PC function. With the recent view on the diversity of the functional role of the cerebellum [97] and the interaction between cerebellum and neocortex [98], the nonlinear dendritic integration of Purkinje cells could potentially contribute to more complex cortical dynamics.
Acknowledgments
This work was partially supported by the National Natural Science Foundation of China under Grant No. 62302016, 62422601, 62176003, and 62088102, and the UK Royal Society Newton Advanced Fellowship Grant No. NAF-R1-191082
Data availability
The code is available at https://github.com/ydtang/Nonlinearity-model.
@@ Acknowledgments
Figure S1: Related to Figure 1. The sublinearity index (%) across the dendritic tree was analyzed in a PC model that incorporates active ion channels. The color-coded morphology (left) illustrates the variation in sublinearity across the dendritic tree. The scatter plots depict the sublinearity index as a function of dendritic diameter (middle) and dendritic distance from the soma (right).
@@
Figure S4: Related to Figure 1. The distributed sublinearity index of each dendrite from three additional PCs. (Left) A rat Purkinje cell (NMO-00891). (Middle) A mouse Purkinje cell (NMO-00864). (Right) A guinea pig Purkinje cell (NMO-00610). The morphology structures were taken from the public archive (www.neuromorpho.org).
Figure S5: Related to Figure3. Histograms show the distribution of characteristics, (left) surface area, (middle) diameter, and (right) length, of 24 branches in the neuron model.
Figure S6: Related to Figure 4 and Figure 6. The Purkinje cell was modeled from guinea pig (NMO_00610). (A). Spiking activity vs. asynchronous input under three spatial stimulation patterns with 50, 100, and 150 synapses. (B). Strong sublinearity suppresses burst. PC installed with 1000 PF AMPA synapses stimulated at 30 Hz Poisson spikes and a CF with 200, 400, and 800 AMPA+NMDA synapses distributed on different sublinear dendrites.
Figure S7: Related to Figure 6. Strong sublinearity suppresses burst. PC installed with 1000 PF AMPA synapses stimulated at 100 Hz Poisson spikes and a CF with 200, 400 and 800 AMPA+NMDA synapses distributed on different sublinear dendrites.
@@
Figure S11: Related to Figure 8. (A) Schematic diagram illustrating the sorting of dendritic branch sublinearity indices (SI) and the implementation of AND operations (cyan) with different numbers of synapses (50, 100, 200, 300) under synchronous and asynchronous input conditions. In general, it is not easy to produce AND operations when the sublinearities of x1 and x2 are both weak. Higher sublinearity indices facilitate AND operations. (B) When executing FBP and dFBP, a schematic diagram of the sublinearity of the four branches.